You expect to recei eapayout from a trust fur d in 5 years The payout wil be for so 200 vou pan to invest the money at an entar-despercent unti the acco anon smso How many years do you have to wat from today? O 1170 years O 800 years 1300 years 0155 years O 975 years
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

20800=12200(1.069)^n

(20800/12200)=(1.069)^n

Taking log on both sides;

log (20800/12200)=n*log (1.069)

n=log (20800/12200)/log (1.069)

=8 years(Approx)

Hence time to wait from today=(8+5)

=13.00 years(Approx).